OpenStudy (anonymous):

what is partial derivative of f(x,y,z,t)=xyz^2tan(yt) what is partial derivative of f(x,y,z,t)=xyz^2tan(yt) @Mathematics

OpenStudy (amistre64):

which partial you want? x,y,z, or t?

OpenStudy (amistre64):

i dunno, you got a funky looking exponent in there that scares me

OpenStudy (amistre64):

its a product so i see a product rule in the future

OpenStudy (amistre64):

but maybe not since the non variables get treated as constant

OpenStudy (amistre64):

F(x,y,z,t)=xyz^2tan(yt) Fx = yz^(2tan(ty)) that is all exponent right?

OpenStudy (anonymous):

with respect to eachj variable jus the z is squared

OpenStudy (amistre64):

\[F(x,y,z,t)=x\ y\ z^2\ tan(yt)\] thats easier then, i hope

OpenStudy (anonymous):

yes the (yt) is what gets me

OpenStudy (amistre64):

Fx = y z^2 tan(yt) Fy = [Cy*tan(yC)]' = C'y*tan(yC)+Cy*tan'(yC) = xz^2 tan(yt) + xyz^2* t sec^2(yt) Fz = 2xyz tan(yt) so far

OpenStudy (anonymous):

it is with respect to y or with respect to t that complex

OpenStudy (amistre64):

Fy means with respect to y Ft might be cleaner tho

OpenStudy (anonymous):

thank you i see its the product rule now

OpenStudy (amistre64):

Ft = [C tan(Ct)]' = C*Csec^2(Ct) = xyz^2 *y sec^2(yt)

OpenStudy (amistre64):

might wanna use different terms for the different constant blocks, but thats the gist of it

OpenStudy (amistre64):

Ft can go to: x y^2 z^2 sec^2(yt)

OpenStudy (anonymous):

r u a teacher/tutor?

OpenStudy (amistre64):

nope, just a person who is trying to remember what i learned about it along time ago :)

OpenStudy (amistre64):

im taking calc3 diffy qs and linear algebra next semester tho

OpenStudy (anonymous):

lol what? retaking them right?

OpenStudy (amistre64):

if i ever have to "retake" a math course, just shoot me ;) nah, im finally going to college to get a degree. Most of this stuff i taught meself.

OpenStudy (amistre64):

well, most of this stuff i taught myself from reading books ... i wish i could have just invented it lol